solve for each system
3x-4y=-6
-5x+3y=-1

Answer by mananth(16946)

3x-4y=-6 .............1
-5x+3y=-1.............2

multiply (1)by 3
Multiply (2) by 4
9x-12y= -18
-20x+12y =-4

Add the two equations
-11x =-22

/-11

x=2

plug value of xin (1)

3x-4y=-6
6-4y=-6
-4y=-6-6
-4y=-12
y=3